OpenStudy (firejay5):

I need help on theses 3 problems. If p(x) = 3x^2 - 2x + 5 and r(x) = x^3 + x + 1, find each value. 30. r(3a) 31. 4p(a) 32. p(a^2)

OpenStudy (tkhunny):

REALLY BIG Hint: p(Frog) = 3(Frog)^2 - 2(Frog) + 5

OpenStudy (firejay5):

that really doesn't help me

OpenStudy (anonymous):

\[r(\spadesuit)=\spadesuit ^3+\spadesuit +1\] replace \(\spadesuit\) by \(3a\)

mathslover (mathslover):

@Firejay5 I will give you a small hint : if P(x) = x^2 - 3 then P(1) = 1^2 - 3 = 1 - 3 = -2

OpenStudy (anonymous):

\[r(3a)=(3a)^3+(3a)+1\]

mathslover (mathslover):

lol @satellite73 nice :)

OpenStudy (anonymous):

actually i liked "frog" better, just couldn't figure out how to make one in latex

OpenStudy (firejay5):

so is (3a)^3 + (3a) + 1 is the answer

OpenStudy (anonymous):

yes, unless you want to change it to \[27a^3+3a+1\]

OpenStudy (firejay5):

what answer does that belong to

OpenStudy (anonymous):

\(r(3a)\)

OpenStudy (tkhunny):

This is a problem in understanding the notation. When you define a function, f(x) = 3x - 5, the terminology f(a) means to substitute the value 'a' everywhere 'x' appears in the definition. f(x) = 3x-5 f(2) = 3(2) - 5 f(2a) = 3(2a) - 5 f(f(x)) = 3(f(x)) - 5 f(Pittsburgh) = 3(Pittsburgh) - 5 f(You must get this idea stuck in your head) = 3(You must get this idea stuck in your head) - 5 Now, let's see your results.

OpenStudy (anonymous):

\[p(x) = 3x^2 - 2x +5 \] means fill in the \(\Box\) \[p(x) = 3\Box^2 - 2\Box +5\]

OpenStudy (firejay5):

which answer is the correct answer

OpenStudy (anonymous):

they are the same, since \((3a)^3=27a^3\)

OpenStudy (firejay5):

@satellite73 what about 32 and then I am done, because I don't get it either

OpenStudy (firejay5):

@Hero Yes???